Please help me solve these problems. I wrote the problems as they are written in my book. Thank You
1.(5,1) is a solution to the equation 2x-3y=7. What does this mean?
2.) Find the missing coordinates
(-3,_)is a solution of y=4-3x

Answer by chessace(471)

1.
(x,y) is standard, so it means x=5 and y=1 will make the equation true, which it does.
2.
x=-3 so y = 4-3(-3) = 4+9 = 13 is the missing 2nd coordinate.